OpenStudy (anonymous):

given the speed of sounf is 340 m/s the frequency is 255 Hz there are 2 speakers (as shown) if a person is walking away towards right from the speakers (red line shows this). What is the distance x where the person no longer hears speakers sound

OpenStudy (stormfire1):

You should probably post the picture

OpenStudy (anonymous):

t1=x1/v, t2=x2/v. x1 and x2 are distances from respective speakers. v=f*wavelength. solve for difference that results in maximum destructive interference.